OpenStudy (anonymous):

1/3(x+6)= 4 Solve for x

OpenStudy (anonymous):

1/3(x+6)=4 1/3*(x+6)/1=4 x+6/3=4(3) x+6=12 Answer: x = 6

OpenStudy (anonymous):

why would you divide by 1?

OpenStudy (anonymous):

why would you divide by 1?

OpenStudy (anonymous):

Oh, it is just easier for me to see and solve it that way:

OpenStudy (anonymous):

1/3 * (x+6)/1 = x+6/3

OpenStudy (anonymous):

i get it now.

OpenStudy (anonymous):

Great. :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You are welcome.